OpenStudy (anonymous):

Identify which lines are perpendicular y=6; y=1/6x ; y=-1/6x; y=-6x

OpenStudy (anonymous):

y=1/6x and y=-6x are perpendicular because 1/6 * (-6) =-1 Similarly y=6 and y=-1/6x are perpendicular because 6 *1/(-6) = -1

OpenStudy (anonymous):

@MeganMckinney

OpenStudy (anonymous):

ah ok, thanks so much!